OpenStudy (gdog):

2. What is a solution of x² + 6x = –5? (1 point)

OpenStudy (mertsj):

Write it in this form: x^2+6x+5=0 and factor it.

OpenStudy (anonymous):

x1= -5 or x2= -1

OpenStudy (anonymous):

We need the quadratic equation to be in the form of: \[ax^2+bx+c=0\] So add 2 to both sides: \[x^2+6x+5=0\] Then factor it.

OpenStudy (gdog):

3. Which of the following is a solution of x² + 14x + 112 = 0? If necessary, round to the nearest hundredth. (1 point)x = –0.24 x = –4.24 4.24 no solution 4. A box shaped like a rectangular prism has a height of 17 in and a volume of 2,720 in³ The length is 4 inches greater than twice the width. What is the width of the box? (1 point)–10 in 8 in 20 in 40 in {1} OK {1} ##LOC[OK]## ##LOC[Cancel]## {1} Yes No